Shift report 20000719 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Jul 19 2000 - 00:00:40 EDT

  • Next message: http account: "Shift report 20000719 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Jul 19 00:00:40 2000, Shift: 16:00-24:00
    Supervisor: BS, with: HI
    Log book#3 pages 27 - 28
    
    Shift Summary:
    ------------------------------------------------------------
    Again RHIC-development all shift, and a power-failure
    in the AGS in addition. They had a few short injections,
    but no ramping and no new beam-schedule.
    
    KO fixed bad TCD in TOFW - all TOFW TDCs lok good, show
    timeouts on pulser trigger.
    
    Due to software-development, the BRAT on the operator-
    account was updated. SuperMon has been rebuilt, and seems
    to be working.
    
    VertexNtuple, a BrNtuple-spinoff has been placed under
    /operator/VertexNtuple and appears to be working fine
    except for some memory-leaks.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Jul 19 2000 - 00:04:06 EDT